The volumes of two HCl solutions A ( 0.5N ) and B ( 0.1N ) to be mixed for preparing 2L of 0.2N HCl are

Grade:12th pass

1 Answers

Vikas TU
7 years ago
From N1V1 = N2V2 we get,
0.5*V1 + 0.1*V2 = 0.2*2
or
5V1 + V2 = 4................(1)
and also obvious,
V1 + V2 = 2...............(2)
Soving the two eqns.we get,
4V1 = 2
v1 =  0.5 L
and
V2 = 1.5 L